### v1

# Ali Azarbahram Shenyu Liu and Gian Paolo Incremona

Abstract
This paper presents a distributed Koopman operator learning framework for modeling unknown nonlinear dynamics using sequential observations from multiple agents. Each agent estimates a local Koopman approximation based on lifted data and collaborates over a communication graph to reach exponential consensus on a consistent distributed approximation. The approach supports distributed computation under asynchronous and resource-constrained sensing. Its performance is demonstrated through simulation results, validating convergence and predictive accuracy under sensing-constrained scenarios and limited communication.
keywords:
Koopman operator, distributed learning, multi-agent systems, nonlinear system identification.
journal: EJC
[…]
[A1]organization=The Department of Electrical Engineering, Chalmers University of Technology,city=Gothenburg, postcode=412 96, country=Sweden
[…]
[A2]organization=The School of Automation, Beijing Institute of Technology,country=China
[…]
[A3]organization=The Dipartimento di Elettronica, Informazione e Bioingegneria, Politecnico di Milano,city=Milan, postcode=20133, country=Italy
1 Introduction
The challenge of modeling and controlling systems with unknown nonlinear dynamics remains central to many engineering and scientific applications. The Koopman operator provides a linear representation of nonlinear systems in a lifted observable space, enabling advanced linear control and prediction techniques Williams et al. (2015). Koopman-based methods have been applied in motion planning Gutow and Rogers (2020), video-based dynamics prediction Comas et al. (2021), and robust quadrotor control Oh et al. (2024). To address model mismatch and uncertainties in practice, Manzoor et al. (2023) surveys hybrid Koopman-based methods with neural networks or robust model predictive control. Moreover, in dynamic environments, Koopman operators have enabled efficient obstacle trajectory prediction and avoidance Lu et al. (2024).
The learning of Koopman operators can be formulated as a data fitting problem, typically solved using least squares algorithms Hansen et al. (2013). However, as system size and complexity increase, particularly in multi-agent and networked settings, centralized Koopman learning becomes impractical due to communication, computation, and scalability limitations. Distributed approaches address these challenges by enabling agents to locally estimate Koopman operators using partial observations and limited communication, thereby improving scalability, preserving privacy, and supporting dynamic network conditions. Implementing such distributed learning requires solving the least squares problem in a distributed manner. Toward this goal, recent advances in distributed algorithms, particularly those in Wang et al. (2019b); Liu et al. (2019); Yang et al. (2020); Huang et al. (2022), achieve exponential convergence guarantees. For broader context, the surveys Wang et al. (2019a); Zheng and Liu (2022) offer comprehensive reviews of related methods.
Recent work has explored various approaches for distributed Koopman operator learning. In Nandanoori et al. (2021), a block-wise Koopman formulation is proposed for multi-agent systems, but the focus remains on block-structured learning under central data access assumptions. The studied approach in Mukherjee et al. (2022) extends this by exploiting graph sparsity and geometric structure, yet their formulation assumes availability of full-row partitions across agents. Deep learning-based methods such as Hao et al. (2024a, b) propose distributed deep Koopman algorithms using neural networks for system identification and control, requiring synchronized neural architectures and backpropagation routines across agents. In Liu et al. (2020), a distributed method is developed under state-based decomposability assumptions to ensure convergence, yet the structure relies on decomposing the dictionary across agents, limiting flexibility in observation patterns. In contrast, our work focuses on the distributed solution of a Frobenius-norm-based Koopman approximation problem with temporal data fragmentation across agents, arising naturally in sensing-constrained scenarios such as unmanned aerial vehicle (UAV) or satellite missions. More precisely, in many practical settings, agents must observe a system sequentially due to bandwidth, sensing, or energy constraints as in satellite missions (e.g., Landsat Wulder et al. (2012)) or multi-UAV sensing and tracking Schwager et al. (2011); Cao et al. (2012). Each agent collects partial temporal snapshots, and information is fused over time to reconstruct an approximation of unknown nonlinear dynamics. The proposed algorithm operates over a generic communication graph and ensures convergence via an augmented-gradient consensus scheme, without requiring raw-data exchange, deep models or parametric dictionaries. The convergence is guaranteed under a unified matrix-theoretic formulation, and validated through rigorous simulations.
The remainder of the paper is structured as follows. The introduction continues with notations and graph definitions; Section II presents the problem formulation and preliminaries. Section III details the proposed method. Section IV illustrates simulation results, and Section V concludes the paper with future directions.
Notations and graph definition
We denote by $\mathbb{N}$ the set of non-negative integers, $\mathbb{R}$ the set of real numbers, $\mathbb{C}$ the set of complex numbers, $\mathbb{R}^{n}$ the $n$-dimensional real space, and $\mathbb{R}^{n\times m}$ the space of $n\times m$ real matrices. In particular, $0_{n\times m}$, (resp. $1_{n\times m}$) denotes the $n\times m$-dimensional zero matrix (resp. all-ones matrix), while $I_{n}$ represents the $n\times n$ identity matrix. When the dimensions are clear from the context, we remove the subindices. A diagonal matrix composed by elements $a_{1},\cdots,a_{p}$ is denoted by $\operatorname{diag}\left(a_{1},\cdots,a_{p}\right)$, and a block diagonal matrix composed by matrices $A_{1},\cdots,A_{p}$ is denoted by $\operatorname{diag}\left(A_{1},\cdots,A_{p}\right)$. The Kronecker product of two matrices $A,B$ is denoted by $A\otimes B$, and the transpose of matrix $A$ is denoted by $A^{\top}$. For square matrix $A\in\mathbb{R}^{n\times n}$, let $\det(A)$ be its determinant, $\operatorname{tr}(A)$ be its trace, and $\Lambda(A)$ be its spectrum (i.e., set of all eigenvalues). For $\lambda\in\mathbb{C}$, let $\operatorname{Re}(\lambda),\,\operatorname{Im}(\lambda)$ be its real and imaginary parts, respectively. For any vector $x\in\mathbb{R}^{n}$, let $\|x\|$ be its 2-norm. For any $A\in\mathbb{R}^{m\times n}$, let $\|A\|_{F}:=\operatorname{tr}(A^{\top}A)$ be its Frobenius norm.
An undirected graph $G=(\mathcal{P},\mathcal{E})$ consists of the vertex set $\mathcal{P}:=\{1,\ldots,p\}$, $p\in\mathbb{N}$, and the edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{P}$, such that $(i,j)\in\mathcal{E}$ if and only if $(j,i)\in\mathcal{E}$. A path is a sequence of vertices connected by edges, and the graph $G$ is connected if there is a path between any pair of vertices. For any $i\in\mathcal{P}$, the set of neighbors of $i$ is $N(i):=\{j\in\mathcal{P}:(i,j)\in\mathcal{E}\}$. A Laplacian matrix $L=[L_{ij}]\in\mathbb{R}^{p\times p}$ of a graph $G$ is given by
$L_{ij}=\begin{cases}-1,&\mbox{if }(i,j)\in\mathcal{E},i\neq j,\\ 0,&\mbox{if }(i,j)\not\in\mathcal{E},i\neq j,\\ -\sum_{k\neq i}L_{ik},&\mbox{if }i=j.\end{cases}$
2 Preliminaries and Problem Statement
The Koopman operator framework enables the study of nonlinear dynamical systems using linear operators acting on functions. Consider a discrete-time nonlinear system of the form
$x_{k+1}$ $=f(x_{k}),\quad x_{k}\in\mathcal{M}\subseteq\mathbb{R}^{q},$ (1)
where $f:\mathcal{M}\to\mathcal{M}$ is a nonlinear map and $q$ is the state dimension. Rather than studying the trajectory of the state $x_{k}$ directly, we analyze the evolution of observables $\psi:\mathcal{M}\to\mathbb{R}$ that belong to a Hilbert space $\mathcal{H}$ of scalar-valued functions. The infinite-dimensional Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ is defined as
$(\mathcal{K}\psi)(x)=\psi(f(x)),\quad\forall\psi\in\mathcal{H},\;x\in\mathcal{M}.$ (2)
Although the system dynamics $f$ is nonlinear, the Koopman operator $\mathcal{K}$ is linear in the space of observables. We select $n$ scalar observables and assemble them into a vector-valued function:
$\Psi(x)$ $=[\psi_{1}(x),\psi_{2}(x),\dots,\psi_{n}(x)]^{\top}\in\mathbb{R}^{n},$ (3)
where $\Psi:\mathcal{M}\to\mathbb{R}^{n}$ maps each state $x\in\mathcal{M}$ to an $n$-dimensional feature space Williams et al. (2015). We then collect $N$ data samples $\{x_{k},f(x_{k})\}_{k=1}^{N}$ and define the data matrices
$X$ $=[\Psi(x_{1}),\Psi(x_{2}),\dots,\Psi(x_{N})]\in\mathbb{R}^{n\times N},$ (4)
$Y$ $=[\Psi(f(x_{1})),\Psi(f(x_{2})),\dots,\Psi(f(x_{N}))]\in\mathbb{R}^{n\times N}.$ (5)
To facilitate computation, the Extended Dynamic Mode Decomposition (EDMD) method Williams et al. (2015) projects the action of $\mathcal{K}$ onto the span of the selected observables. This results in a matrix $K\in\mathbb{R}^{n\times n}$ that approximates the Koopman operator over the chosen feature space and governs the evolution of lifted states in the observable space. Hence, the finite-dimensional approximation of the Koopman operator $\mathcal{K}$ via EDMD seeks a matrix $K\in\mathbb{R}^{n\times n}$ that best satisfies
$Y$ $\approx KX.$ (6)
This leads to the following least-squares optimization using the Frobenius norm:
$K^{*}$ $=\arg\min_{K\in\mathbb{R}^{n\times n}}\left\|Y-KX\right\|_{F}^{2}.$ (7)
In a distributed setting, each agent $i\in\{1,\ldots,p\}$ collects its own segment of the system observations, resulting in local data matrices $Y_{i}\in\mathbb{R}^{n\times m_{i}}$ and $X_{i}\in\mathbb{R}^{n\times m_{i}}$, where $m_{i}$ denotes the number of temporal snapshots available to agent $i$. These data blocks represent lifted observables evaluated at the agent’s local state transitions. Such partitioning naturally arises in sequential observation frameworks, where agents take turns sensing the system over time, or when access to data is constrained by communication, privacy, or coverage limitations. To construct a distributed Koopman approximation, the locally collected data are aggregated to form the global matrices
$Y=\begin{bmatrix}Y_{1}&Y_{2}&\cdots&Y_{p}\end{bmatrix},\quad X=\begin{bmatrix}X_{1}&X_{2}&\cdots&X_{p}\end{bmatrix}$,
with $Y,\,X\in\mathbb{R}^{n\times N}$ and $\sum_{i=1}^{p}m_{i}=N$. This structure enables decentralized computation while preserving the full representation of the system’s evolution.
From the problem formulation, the $i$-th agent only knows $Y_{i},X_{i}$. Because $\|Y-KX\|_{F}^{2}~=\sum_{i=1}^{p}\|Y_{i}-KX_{i}\|_{F}^{2}$, we could study the distributed problem that each agent aims to find a solution $K_{i}\in\mathbb{R}^{n\times n}$ for the local problem $\min_{K_{i}}\|Y_{i}-K_{i}X_{i}\|_{F}$, subject to the constraint that all $K_{i}$ must be equal. Furthermore, denote
$\mathbf{K}:$ $=\begin{bmatrix}K_{1}&K_{2}&\cdots&K_{p}\end{bmatrix}\in\mathbb{R}^{n\times np},$ (8)
$\bm{X}:$ $=\operatorname{diag}(X_{1},X_{2},\cdots,X_{p})\in\mathbb{R}^{np\times N},$ (9)
$\mathbf{L}:$ $=L\otimes I_{n}\in\mathbb{R}^{np\times np},$ (10)
where recall $L$ is the Laplacian matrix of the communication graph $G$. We further have that $\|Y-KX\|_{F}^{2}=\sum_{i=1}^{p}\|Y_{i}-KX_{i}\|_{F}^{2}=\|Y-\mathbf{K}\bm{X}\|_{F}^{2}$, subject to the constraint $K=K_{1}=K_{2}=\cdots=K_{p}$; that is, $\mathbf{K}=1_{1\times p}\otimes K$, which can be equivalently expressed as $\mathbf{K}\mathbf{L}=0$ since the communication graph $G$ is connected. Hence, to solve (7) by a distributed method in this scenario, we can equivalently study the optimization problem
$\min_{\mathbf{K}\in\mathbb{R}^{n\times np}}$ $\frac{1}{2}\|Y-\mathbf{K}\bm{X}\|_{F}^{2},$ (11a)
$\mbox{subject to }\mathbf{K}\mathbf{L}=0.$ (11b)
3 Main Results
We propose to solve the problem (7) by a discrete-time distributed algorithm. Essentially, the agents alternate between communication and computation, such that they run the following updating law
$K_{i}^{+}$ $=K_{i}-\alpha\Big((K_{i}X_{i}-Y_{i})X_{i}^{\top}$
$~~~~~~~~~~~~~~~~~+k_{P}\sum_{j\in N(i)}(K_{i}-K_{j})+k_{I}R_{i}\Big),$ (12a)
$R_{i}^{+}$ $=R_{i}+\alpha\sum_{j\in N(i)}(K_{i}-K_{j}).$ (12b)
Here, $\alpha>0$ is a fixed step-size, $k_{p},\,k_{I}>0$ are some tunable gains, $K_{i}\in\mathbb{R}^{n\times n}$ is the $i$-th agent’s individual guess of the optimum of (7), $R_{i}\in\mathbb{R}^{n\times n}$ is another internal augmented state variable owned by the $i$-th agent for facilitating the convergence. The distributed algorithm for Koopman operator learning is summarized in Algorithm 1. The update law is similar to the PI-control studied in Yang et al. (2019), and we provide some intuitive interpretations which help understand Algorithm 1.
  • The term $(Y_{i}-K_{i}X_{i})X_{i}^{\top}$ in (12a) is the gradient of the local cost $\frac{1}{2}\|Y_{i}-K_{i}X_{i}\|_{F}^{2}$, indicating that each agent minimizes its own Koopman approximation error.
  • The term $\sum_{j\in N(i)}(K_{i}-K_{j})$ in (12b) acts as a proportional diffusion term, promoting consensus among neighboring agents.
  • Given $R_{i}(0)=0$, the update in (12b) shows that $R_{i}$ accumulates the diffusion term over time, functioning as an integral feedback that enforces consensus across agents.
The convergence of Algorithm 1 is related to the eigenvalues of the following matrix
$\mathbf{M}:=\begin{bmatrix}-\bm{X}\bm{X}^{\top}-k_{P}\mathbf{L}&\mathbf{L}\\ -k_{I}I_{np}&0\end{bmatrix}\in\mathbb{R}^{2np\times 2np},$ (13)
where recall $\bm{X},\mathbf{L}$ are defined in (9), (10). Note that $\mathbf{M}$ depends on the gains $k_{P},\,k_{I}$. The following theorem provides the design criteria of these parameters and the step-size $\alpha$ which guarantees the convergence of Algorithm 1, and relates the convergence rate to these parameters.
Theorem 1.
Suppose the communication graph $G$ is undirect and connected. For any gains $k_{P},k_{I}>0$, there exists
[…]
such that as long as the step size $\alpha<\alpha_{\max}$, the $K_{i}$ components of the algorithm will reach consensus and converge to an optimal solution of the problem (7). Furthermore, the convergence is exponential with rate $\rho>\rho_{\max}$, where
[…]
The proof of Theorem 1 is provided in the Appendix. We note that the upper bound $\alpha_{\max}$ in (14) on the step size must be computed using the matrix $\mathbf{M}$, which depends on the centralized quantities $X,L$. Consequently, $\alpha_{\max}$ cannot be determined a priori in a distributed setting. Our convergence guarantee for Algorithm 1 should be interpreted qualitatively: provided the step size is sufficiently small, the algorithm converges exponentially fast to an optimal solution of (7). Future work will explore the development of distributed algorithms that either incorporate adaptive step sizes or eliminate the need for step-size tuning altogether.
Algorithm 1 Distributed Koopman Operator Learning
[…]
2:Initialize arbitrary $K_{i}(0)\in\mathbb{R}^{n\times n}$, $R_{i}(0)=0_{n\times n}$ for all $i\in\mathcal{P}$.
3:loop for $t=0,1,\ldots,t_{\max}-1$, the $i$-th agent, $i\in\mathcal{P}$
4:  Broadcast $K_{i}(t)$ to its neighbors.
5:  Compute $K_{i}(t+1),R_{i}(t+1)$ according to (12).
6:end loop
[…]
Figure 1: Spectral comparison between $K_{\mathrm{ave}}$ and $K^{*}$.
Figure 2: Elementwise absolute difference between $K_{\mathrm{ave}}$ and $K^{*}$.
4 Simulation results
We test the distributed Koopman learning method on a $20\times 20$ grid capturing a nonlinear process (e.g., pedestrian density) via top-down normalized intensity maps in $[0,1]$. Three camera agents each gather $m_{i}=3$ local snapshots ($N=9$ total), and collaborate without centralized fusion. Each $20\times 20$ frame is vectorized to $n=400$ for Koopman lifting, with agents solving (7) in a distributed manner. Communication follows a circular undirected graph $G$ encoded by Laplacian $L$, reflecting decentralized conditions. Each agent runs $1000$ iterations of the distributed update, using $k_{P}=150$, $k_{I}=75$, and $\alpha=0.5\alpha_{\max}$ with $\alpha_{\max}=0.03$ from (14). This ensures exponential convergence with $\rho_{\max}=0.92$ as per (15).
Figure 3: Prediction error of distributed Koopman operators.
Figure 4: Heatmap of distributed Koopman prediction error over time.
We analyze the spectral similarity between centralized and distributed Koopman operators. Let $K^{*}$ be the centralized solution to (7) using $(X,Y)$, and $K_{i}$ the local estimate from agent $i$. Define the average distributed operator as $K_{\mathrm{ave}}:=1/p\sum_{i=1}^{p}K_{i}$. Figure 1 shows that $K^{*}$ and $K_{\mathrm{ave}}$ share dominant eigenvalues near the unit circle, indicating accurate capture of persistent dynamics. Eigenvalues near the origin in $K_{\mathrm{ave}}$ reflect suppressed modes from limited observability and averaging, contributing to robustness. Figure 2 shows entrywise differences $|K_{\mathrm{ave}}-K^{*}|$. Most values are near zero, with block structures matching spatially-local observables on the $20\times 20$ grid. Minor deviations (around $10^{-4}$) likely stem from observability gaps or communication limits, demonstrating the method’s resilience and offering insight for future refinements. To assess accuracy, we compute $1/p\sum_{i=1}^{p}\left\|Y-K_{i}X\right\|_{F}$, measuring how well each distributed Koopman operator $K_{i}$ generalizes to the full dataset $(X,Y)$. Despite being trained on local data, agents achieve low reconstruction error over time, as shown in Figure 3, indicating alignment with global measurements and inter-agent consensus. To evaluate the predictions, we simulate the predictions and compare it to ground truth. Figure 4 shows this absolute error over time and space. Errors stay low ($\sim 10^{-2}$) and structured, confirming that $K_{\mathrm{ave}}$ enables accurate, stable multi-step predictions across the spatial domain, even under decentralization. While our results confirm convergence to the centralized least-squares solution, we also emphasize that the low multi-step prediction error (Figure 4) empirically supports the Koopman model’s ability to approximate the underlying dynamics and not just fit the training data. This confirms the model’s practical forecasting capability beyond mere algebraic consistency.
Remark 1.
In our simulations, we employ direct vectorization of the $20\times 20$ grid as observables. This choice is natural for representing spatial phenomena such as pedestrian density and preserves locality in the lifted space. It enables interpretable modeling while maintaining computational tractability. Given our focus on distributed Koopman operator learning under distributed data access, this observable structure offers a practical and consistent approach for operator estimation across agents. We note that all agents observe temporally partitioned data from the same environment using a shared lifting function, rather than agent-specific observables.
Remark 2.
Data length influences Koopman approximation by determining coverage of the system’s state space. In our case, $p=3$ agents each gather $m_{i}=3$ snapshots ($N=9$ total), which sufficed for approximately accurate multi-step prediction (Figures 3 and 4). This short horizon reflects realistic constraints (e.g., limited bandwidth or battery). Despite this, the learned $K_{i}$ generalize well and capture actual dynamics approximately. The Koopman operator is estimated via least-squares over lifted observables. Its spectral similarity to the centralized solution (Figure 1) and strong prediction performance further confirm the approximation quality. Moreover, the lifted dimension $n=400$ provides sufficient expressive power to capture dominant spatial modes, while maintaining computational feasibility for distributed implementation.
We benchmark centralized and distributed implementations in MATLAB (Intel i7, 16GB RAM) for a $20\times 20$ grid ($n=400$). Centralized learning, based on $K=YX^{\dagger}$ (with $X^{\dagger}$ denoting the Moore–Penrose pseudoinverse), completes in $\sim$15.2 ms. Distributed learning takes $\sim$0.3 ms for local computation and $\sim$0.1 ms for communication per iteration (averaged over 1000 steps), totaling $\sim$400 ms. Although centralized learning appears faster for small datasets, it requires full data aggregation and becomes more computationally expensive as the number of snapshots increases. Moreover due to energy, bandwidth, or field-of-view constraints, continuous high-resolution data collection from a single agent is often physically infeasible in real-world scenarios. To contextualize our contribution, we compare against the distributed Koopman framework proposed in Nandanoori et al. (2021). Under the same simulation setup, we implemented Nandanoori et al. (2021) with a minor adaptation to support shared observations, allowing comparison under identical conditions. While their approach suits modular systems with well-defined inter-agent couplings, it lacks consistency guarantees in shared, temporally fragmented observations. Our method improves multi-step prediction by $\sim 21\%$, due to integral feedback in Algorithm 1.
Remark 3.
The convergence of Algorithm 1 is driven by the spectrum of $\mathbf{M}$ in (13), shaped by $k_{P}$, $k_{I}$:
  • $k_{P}$: Enhances consensus by penalizing disagreement. Too high may suppress local fitting.
  • $k_{I}$: Reinforces long-term agreement but risks instability unless $\alpha$ is small.
  • $\alpha$: Must satisfy $\alpha<\alpha_{\max}$ from (14). Larger $k_{P}$, $k_{I}$ reduce the spectral radius of $\mathbf{M}$, increasing $\alpha_{\max}$ and improving the convergence rate governed by $\rho_{\max}$ in (15). Set $\alpha=\theta\alpha_{\max}$ with $\theta\in(0.3,0.6)$.
  • Practice: Use moderate gains, compute $\alpha_{\max}$, and set $\alpha=0.5\alpha_{\max}$. To accelerate, increase $k_{P}$; for stability, decrease $\alpha$ or $k_{I}$. Choose $t_{\max}$ based on $\rho_{\max}$.
Remark 4.
In practice, each agent collects local data over a brief sequence of time steps, after which Algorithm 1 is executed. This interpretation of sequential observations aligns with physical constraints such as frame-rate or communication delays in agent sensing. Figure 3 shows that the distributed Koopman operators $K_{i}$ achieve low reconstruction error after around 500 iterations, which occur on millisecond time scales that is significantly faster than the few-second intervals at which video snapshots (e.g., of pedestrian motion) typically arrive. This separation ensures convergence is feasible before new data becomes available. As confirmed in Figure 4, the average operator $K_{\mathrm{ave}}$ provides accurate multi-step predictions even under decentralization, enabling online forecasting while new data is being acquired. For example, as three agents collect the next $N=9$ snapshots (each with $m_{i}=3$), the current $K_{i}$’s can be used for short-term prediction. Once the data is available, the algorithm is rerun to refine the models, making the framework suitable for real-time deployment in applications such as crowd monitoring.
Remark 5.
While each agent maintains a Koopman matrix $K_{i}\in\mathbb{R}^{n\times n}$ which scales quadratically with $n$, this is practical in scenarios such as satellite or UAV-based sensing, where agents acquire time-varying observations under bandwidth or energy constraints. The focus in our proposed appraoch is on estimating consistent Koopman models under distributed and resource-limited conditions, rather than globally approximating nonlinear dynamics mitigating the dimensionality issue noted in Zhang and Zuazua (2023). Future extensions may incorporate scalable lifting strategies (e.g., sparse or low-rank forms) to address larger-scale problems.
Remark 6.
Algorithm 1 converges exponentially with rate governed by $\rho_{\max}$ in (15). As shown in Figure 3, approximately an accurate reconstruction is achieved after $\sim$500 iterations. In online settings, where time and resources are limited, sub-optimal iterates can still be used for short-term prediction. Figure 4 shows that even intermediate solutions yield reliable forecasts. As new data arrives, the algorithm can resume from the latest iterate, enabling continual refinement. This supports real-time deployment, where learning proceeds alongside sensing and control.
5 Conclusion
This work proposed a distributed Koopman learning algorithm for modeling unknown nonlinear dynamics from sequential agent-level observations. Local Koopman models are constructed from lifted temporal data and refined via consensus to achieve consistent overall representations under distributed constraints. Simulation results demonstrate convergence and approximate model reconstruction using the proposed distributed learning method. In future, the learned Koopman operator can support real-time prediction of unknown motions, enabling proactive coordination and decision-making for dynamic obstacle avoidance in multi-agent scenarios.
Appendix A
The proof of Theorem 1 relies on the following Lemma.
Lemma 1.
Consider a matrix
$\tilde{\mathbf{M}}=\begin{bmatrix}-\bm{X}\bm{X}^{\top}-k_{P}\mathbf{L}&\sqrt{k_{I}}\mathbf{L}^{\frac{1}{2}}\\ -\sqrt{k_{I}}\mathbf{L}^{\frac{1}{2}}&0\end{bmatrix}$. (16)
[…]
Proof.
Since the matrix $\mathbf{L}^{\frac{1}{2}}$ and $I$ commute, we use Schur complement and determinantal formula Horn and Johnson (2012, Chapter 0.8.5) to conclude
[…]
$=\det(s(sI+\bm{X}\bm{X}^{\top}+k_{P}\mathbf{L})+k_{I}\mathbf{L})$
[…]
Proof of Theorem 1.
Similar to the $\mathbf{K}$ defined in (8), define
$\mathbf{R}:=\begin{bmatrix}R_{1}&R_{2}&\cdots&R_{p}\end{bmatrix}\in\mathbb{R}^{n\times np}.$
The update law (12b) and the initial state $R_{i}(0)=0_{n\times n}$ imply the existence of $\tilde{\mathbf{R}}(k)\in\mathbb{R}^{n\times np}$ such that $\mathbf{R}(k)=\frac{1}{\sqrt{k_{I}}}\tilde{\mathbf{R}}(k)\mathbf{L}^{\frac{1}{2}}$ for all $k\in\mathbb{N}$. Therefore, the update law (12) can be rewritten as
$\begin{bmatrix}\mathbf{K}^{+}&\tilde{\mathbf{R}}^{+}\end{bmatrix}=\begin{bmatrix}\mathbf{K}&\tilde{\mathbf{R}}\end{bmatrix}(I+\alpha\tilde{\mathbf{M}})+\alpha\begin{bmatrix}Y\bm{X}^{\top}&0\end{bmatrix}$, (17)
where we recall that the matrices $\bm{X},\tilde{\mathbf{M}}$ are defined in (9), (16). Meanwhile, it follows from the KKT condition Boyd and Vandenberghe (2004) that for any optimal solution $\mathbf{K}^{*}\in\mathbb{R}^{n\times np}$ of the constrained problem (11), there exists $\Lambda^{*}\in\mathbb{R}^{n\times np}$ such that
$-(Y-\mathbf{K}^{*}\bm{X})\bm{X}^{\top}+\Lambda^{*}\mathbf{L}$ $=0,$
$\mathbf{K}^{*}\mathbf{L}$ $=0.$
These equations can be compactly written as
$\begin{bmatrix}\mathbf{K}^{*}&-\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}\tilde{\mathbf{M}}+\begin{bmatrix}Y\bm{X}^{\top}&0\end{bmatrix}=0.$ (19)
For each $k\in\mathbb{N}$, define $\mathbf{E}(k):=\begin{bmatrix}\mathbf{E}_{K}(k)&\mathbf{E}_{R}(k)\end{bmatrix}$, where
$\mathbf{E}_{K}(k):=\mathbf{K}(k)-\mathbf{K}^{*},\quad{\color[rgb]{0,0,0}\mathbf{E}_{R}(k):=\tilde{\mathbf{R}}(k)+\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}.}$
It follows from (17) and (19) that
$\mathbf{E}^{+}$ $=\begin{bmatrix}\mathbf{K}^{+}&\tilde{\mathbf{R}}^{+}\end{bmatrix}-\begin{bmatrix}\mathbf{K}^{*}&-\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}$
$=\begin{bmatrix}\mathbf{K}&\tilde{\mathbf{R}}\end{bmatrix}(I+\alpha\tilde{\mathbf{M}})+\alpha\begin{bmatrix}Y\bm{X}^{\top}&0\end{bmatrix}\quad-\begin{bmatrix}\mathbf{K}^{*}&\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}$
$=\begin{bmatrix}\mathbf{K}&\tilde{\mathbf{R}}\end{bmatrix}(I+\alpha\tilde{\mathbf{M}})-\alpha\begin{bmatrix}\mathbf{K}^{*}&-\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}\tilde{\mathbf{M}}\quad-\begin{bmatrix}\mathbf{K}^{*}&\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}$
[…]
which is a matrix-valued linear time-invariant system. Hence, we only need to analyze the stability properties of the matrix $I+\alpha\tilde{\mathbf{M}}$. For any eigenvalue $\lambda^{\prime}$ of $I+\alpha\tilde{\mathbf{M}}$, $\lambda^{\prime}=1+\alpha\operatorname{Re}(\lambda)+\alpha\operatorname{Im}(\lambda)j$, where $\lambda$ is an eigenvalue of $\tilde{\mathbf{M}}$. Hence
$|\lambda^{\prime}|^{2}=(1+\alpha\operatorname{Re}(\lambda))^{2}+\alpha^{2}\operatorname{Im}(\lambda)^{2}=1+2\alpha\operatorname{Re}(\lambda)+\alpha^{2}|\lambda|^{2}.$
Meanwhile, by Liu (2024, Lemma 7), $\tilde{\mathbf{M}}$ is semi-Hurwitz; that is, all its eigenvalues either have negative real parts, or are 0 and non-defective (algebraic multiplicity equals to geometric multiplicity) Additionally, because of Lemma 1 and the assumption $\alpha<\alpha_{\max}$, where $\alpha_{\max}$ is defined in (14), we either have $\lambda=0$ and non-defective so that $\lambda^{\prime}=1$ and non-defective, or $\operatorname{Re}(\lambda)<0$ and $|\lambda^{\prime}|<1$. This implies that $I+\alpha\tilde{\mathbf{M}}$ has a Jordan decomposition of the form
$I+\alpha\tilde{\mathbf{M}}=\begin{bmatrix}V_{1}&V_{0}\end{bmatrix}\begin{bmatrix}J_{1}\\ &I\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{0}\end{bmatrix}=:VJU,$ (21)
such that $UV=I$, $J_{1}\in\mathbb{R}^{n^{\prime}\times n^{\prime}}$ for some $n^{\prime}<2np$ is a block-diagonal matrix of Jordan blocks. Moreover, $J_{1}$ is Schur (all eigenvalues are in the open unit disk) and the spectral radius of $J_{1}$ is $\rho_{\max}$, defined in (15). Since $\rho>\rho_{\max}$, $\frac{1}{\rho}J_{1}$ is still Schur and hence by the discrete-time Lyapunov equation, there exists a symmetric positive definite matrix $P\in\mathbb{R}^{n^{\prime}\times n^{\prime}}$ such that
$\frac{1}{\rho^{2}}J_{1}PJ_{1}^{\top}-P\prec 0,$ (22)
where $\prec 0$ means the matrix is negative definite. For each $k\in\mathbb{N}$, denote $\mathbf{F}_{1}(k):=\mathbf{E}(k)V_{1},\mathbf{F}_{0}(k):=\mathbf{E}(k)V_{0}$.
It follows from (20) and (21) that
$\mathbf{F}_{0}^{+}$ $=\mathbf{E}^{+}V_{0}=\mathbf{E}\begin{bmatrix}V_{1}&V_{0}\end{bmatrix}\begin{bmatrix}J_{1}\\ &I\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{0}\end{bmatrix}V_{0}$
$=\mathbf{E}V_{0}=\mathbf{F}_{0}.$
Hence $\mathbf{F}(k)=\mathbf{F}(0)$. Similarly,
$\mathbf{F}_{1}^{+}$ $=\mathbf{E}^{+}V_{1}=\mathbf{E}\begin{bmatrix}V_{1}&V_{0}\end{bmatrix}\begin{bmatrix}J_{1}\\ &I\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{0}\end{bmatrix}V_{1}$
$=\mathbf{E}V_{1}J_{1}=\mathbf{F}_{1}J_{1}.$
Then, it further follows from (22) that
$\|\mathbf{F}_{1}^{+}P^{\frac{1}{2}}\|_{F}^{2}$ $=\operatorname{tr}(\mathbf{F}_{1}^{+}P(\mathbf{F}_{1}^{+})^{\top})=\operatorname{tr}(\mathbf{F}_{1}J_{1}PJ_{1}^{\top}\mathbf{F}_{1}^{\top})$
$\leq\rho^{2}\operatorname{tr}(\mathbf{F}_{1}P\mathbf{F}_{1}^{\top})=\rho^{2}\|\mathbf{F}_{1}P^{\frac{1}{2}}\|_{F}^{2},$
where we have used the fact that the trace of positive semi-definite matrix is non-negative for the inequality above. Hence, $\|\mathbf{F}_{1}(k)P^{\frac{1}{2}}\|_{F}\leq\rho^{k}\|\mathbf{F}_{1}(0)P^{\frac{1}{2}}\|_{F}$. In other words, $\mathbf{F}_{1}(k)$ converges to 0 exponentially with rate $\rho$. Because $\mathbf{E}=\begin{bmatrix}\mathbf{F}_{1}&\mathbf{F}_{0}\end{bmatrix}U=\mathbf{F}_{1}U_{1}+\mathbf{F}_{0}U_{0}=\mathbf{F}_{1}U_{1}+\mathbf{E}V_{0}U_{0}$, we conclude that $\mathbf{E}(k)$ converges to $\mathbf{E}_{\infty}:=\mathbf{E}(0)V_{0}U_{0}$ exponentially with rate $\rho$. In addition,
$\mathbf{E}_{\infty}(I+\alpha\tilde{\mathbf{M}})=\mathbf{E}(0)V_{0}U_{0}\begin{bmatrix}V_{1}&V_{0}\end{bmatrix}\begin{bmatrix}J_{1}\\ &I\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{0}\end{bmatrix}\\ =\mathbf{E}(0)\begin{bmatrix}0&V_{0}\end{bmatrix}\begin{bmatrix}J_{1}\\ &I\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{0}\end{bmatrix}=\mathbf{E}(0)V_{0}U_{0}=\mathbf{E}_{\infty},$
which implies
$\mathbf{E}_{\infty}\tilde{\mathbf{M}}=0.$ (23)
Recall the definition of $\mathbf{E}$. The solution of (12), $\begin{bmatrix}\mathbf{K}(k)&\tilde{\mathbf{R}}(k)\end{bmatrix}$, converges to
$\begin{bmatrix}\mathbf{K}_{\infty}&\mathbf{R}_{\infty}\end{bmatrix}:=\mathbf{E}_{\infty}+\begin{bmatrix}\mathbf{K}^{*}&-\frac{1}{\sqrt{k_{I}}}\Lambda^{*}\mathbf{L}^{\frac{1}{2}}\end{bmatrix}$.
It then follows from (19) and (23) that
$\begin{bmatrix}\mathbf{K}_{\infty}&\mathbf{R}_{\infty}\end{bmatrix}\tilde{\mathbf{M}}+\begin{bmatrix}Y\bm{X}^{\top}&0\end{bmatrix}=0.$
In other words, $\mathbf{K}_{\infty},\mathbf{R}_{\infty}$ also satisfy the KKT condition of optimality for (11). Hence, $\mathbf{K}_{\infty}$ is an optimal solution for the problem (11), and $\mathbf{K}_{\infty}=\begin{bmatrix}K_{\infty}&K_{\infty}&\cdots&K_{\infty}\end{bmatrix}$, where $K_{\infty}\in\mathbb{R}^{n\times n}$ is an optimal solution for the problem (7). This completes the proof. ∎